# Entanglement transfer during quantum frequency conversion in gas-filled hollow-core fibers

### v1

@@ III Equations of motion
using the formulas
$[\sigma_{z},\sigma^{+}]$ $=$ $2\sigma^{+},$
$[\sigma_{z},\sigma^{-}]$ $=$ $-2\sigma^{-},$
$[\sigma^{+},\sigma^{-}]$ $=$ $\sigma_{z}.$
By defining $w=\rho_{1,1}-\rho_{0,0}$, we write the following EOM,
where we have not written the equation for $\dot{\rho}_{10}$ because it is just the complex conjugate of that for $\dot{\rho}_{0,1}$. The differences we find with the previous results lie in the definition of the electric field. While we have defined $\mathcal{E}(z,t)=\mathcal{E}_{P}(z,t)+\mathcal{E}_{S}(z,t)$, previous works defined $\mathcal{E}(z,t)=(\mathcal{E}_{P}(z,t)+\mathcal{E}_{S}(z,t))/2$, and thus the extra factor of $1/4$. Furthermore, while we have chosen the plain wave solution of the electric fields as
previous works have chosen
and this is why we find $E_{P}^{*}E_{S}$ instead of $E_{P}E_{S}^{*}$.
@@ IV Spin Coherent State
Notice that, in resonance conditions, $\Omega-\omega_{P}+\omega_{S}=0$, and there is no explicit time dependence in the Hamiltonian. Finally, we perform a semiclassical approximation on the pump and Stokes modes, $a_{l}\rightarrow\alpha_{l}$, and obtain
@@ V Concurrence
After coherence has been generated in the molecule ensemble, we will send one mode of an entangled state through the fiber to change its frequency. Unlike coherence modulation, this is a thresholdless process which can be applied to single photons. We consider $(|0,0,0\rangle+|1,1,0\rangle)/\sqrt{2}$ as the initial state, where the states in tensor product, from left to right, indicate the idler mode, kept outside of the fiber, the mixing mode, which goes initially through the fiber, and the frequency-converted mode, initially in a vacuum state. While the phase-difference between the original and the frequency-converted signals has to be equal to the phase of the molecular coherence wave, the jump in frequency is given by the Raman shift. Therefore, even though this is not a Raman process, we can use the same interaction Hamiltonian to describe it. Furthermore, we will focus on the interaction between the pump and the anti-Stokes, since we are looking at frequency up-conversion. Then, we start from the Hamiltonian
@@
Under phase matching and resonance conditions, we have $\beta_{U}-\beta_{M}=\Delta\beta$ and $\omega_{U}-\omega_{M}=\Omega$, respectively. Assuming these are satisfied, this Hamiltonian simplifies to
@@
As we can see, this resembles a beamsplitter operator. Now, we need to obtain the action of this Hamiltonian onto the initial state of the system, which will be $(|0,0,0\rangle+|1,1,0\rangle)/\sqrt{2}$. But first, we need to transform this state into the interaction picture,
@@
We obtain the idler-mixing density matrix by tracing out the up-converted mode,
@@
and the idler-up-converted density matrix is obtained by tracing out the mixing mode,
@@
$\mathcal{C}=\max\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\},$ (S98)
@@
and hence, the concurrence is given by $\mathcal{C}_{I-M}=\max\{0,2|w|\}$. In the case of $\rho_{I,U}$, we have
@@
finding the concurrence $\mathcal{C}_{I-U}=\max\{0,2|y|\}$. Therefore, the concurrences can be computed as
$\mathcal{C}_{I,M}$ $=$ $\Big{|}\cos G_{U}|\xi|t\Big{|},$ (S107)
$\mathcal{C}_{I,U}$ $=$ $\Big{|}\sin G_{U}|\xi|t\Big{|}.$ (S108)
@@
$\mathcal{E}_{\rho}=1-\frac{1}{2}\left[\left(1+\sqrt{1-\mathcal{C}^{2}}\right)\log_{2}\left(1+\sqrt{1-\mathcal{C}^{2}}\right)+\left(1-\sqrt{1-\mathcal{C}^{2}}\right)\log_{2}\left(1-\sqrt{1-\mathcal{C}^{2}}\right)\right].$ (S109)
@@ VI Concurrence without semiclassical approximation for spins
In this section, we will look into the semiclassical approximation for spin operators that we performed before computing the concurrence between the idler-mixing and the idler-up-converted modes. This approximation is based on the same one for bosonic states, in which the system is assumed to be in a coherent state $|\alpha\rangle$, and thus we replace $a\rightarrow\alpha$. This assumes that the system remains in a coherent state, which is what we are considering here with the molecules; after these are in a coherent state, we are looking at the dynamics after introducing a single-excitation into the system. Let us derive here the idler-mixing and idler-up-converted states without the semiclassical approximation of the spin operators. We start from the state
@@
in the Schrödinger picture, which describes the idler, mixing, up-converted, and molecules modes, in that order. We want to obtain the evolution of this state under our Hamiltonian
@@
where we have already implemented phase-matching condition $\beta_{U}-\beta_{M}=\Delta\beta$. Again, we now move to an interaction picture with respect to $\hbar\Omega J_{z}+\hbar\omega_{M}a_{M}^{\dagger}a_{M}+\hbar\omega_{U}a_{U}^{\dagger}a_{U}$, where we have
@@
The density matrix we obtain after tracing out the molecules is
@@
By tracing the up-converted mode, we obtain the idler-mixing density matrix,
@@
while the density matrix for the idler and the up-converted modes is obtained by tracing the mixing mode,
@@
$x$ $=$ $\sum_{n=0}^{N}|c_{1}(n)|^{2}=\frac{1}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N}\begin{pmatrix}N\\ n\end{pmatrix}|s|^{2n}\cos^{2}(G_{U}t\sqrt{n(N-n+1)}),$
$w$ $=$ $\sum_{n=0}^{N}c_{1}(n)c_{0}^{*}(n)=\frac{1}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N}\begin{pmatrix}N\\ n\end{pmatrix}|s|^{2n}\cos(G_{U}t\sqrt{n(N-n+1)}),$
$y$ $=$ $\sum_{n=0}^{N-1}c_{2}(n+1)c_{0}^{*}(n)=-ie^{-i(\Omega t+\Delta\beta z)}\frac{s}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N-1}\begin{pmatrix}N\\ n\end{pmatrix}\sqrt{\frac{N-n}{n+1}}|s|^{2n}\sin(G_{U}t\sqrt{(n+1)(N-n)}).$
@@
$w=\frac{1}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N}\begin{pmatrix}N\\ n\end{pmatrix}|s|^{2n}\cos(G_{U}t\sqrt{n(N-n+1)}),$
$w_{\text{sc}}=\frac{1}{2}\cos G_{U}|\xi|t.$
@@
$\frac{G_{U}t}{8}\left\{-NG_{U}t\cos^{2}(2G_{S}\alpha_{P}\alpha_{S}t)\cos G_{U}|\xi|t+\left[\cot(2G_{S}\alpha_{P}\alpha_{S}t)-\tan(2G_{S}\alpha_{P}\alpha_{S}t)\right]\sin G_{U}|\xi|t\right\}.$ (S131)
@@
$\frac{1}{8N}\left\{-G_{U}^{2}t^{2}|\xi|^{2}\left(1-\cot^{2}(2G_{S}\alpha_{P}\alpha_{S}t)\right)^{2}\cos G_{U}|\xi|t+G_{U}t|\xi|\left(\frac{1}{\sin^{2}(2G_{S}\alpha_{P}\alpha_{S}t)}-\frac{3}{\cos^{2}(2G_{S}\alpha_{P}\alpha_{S}t)}\right)\sin G_{U}|\xi|t\right\}.$ (S132)

### v2

@@ III Semiclassical approximation for photon modes
$\rho^{(2)}=\frac{1}{2}\left(\frac{it}{\hbar}\right)^{2}\left[H_{\text{eff}}^{I},\left[H_{\text{eff}}^{I},|\alpha_{P}\rangle\otimes|\alpha_{S}\rangle\otimes|\Psi\rangle\langle\alpha_{P}|\otimes\langle\alpha_{S}|\otimes\langle\Psi|\right]\right]$ (S45)
By taking the partial trace over the bosonic modes, we find
[…]
$+$ $\frac{1}{2}\left(\frac{it}{\hbar}\right)^{2}\hbar^{2}|G_{s}|^{2}\Big{[}|\alpha_{P}|^{2}\left(J^{-}J^{+}|\Psi\rangle\langle\Psi|+|\Psi\rangle\langle\Psi|J^{-}J^{+}-2J^{+}|\Psi\rangle\langle\Psi|J^{-}\right)$
$+$ $|\alpha_{S}|^{2}\left(J^{+}J^{-}|\Psi\rangle\langle\Psi|+|\Psi\rangle\langle\Psi|J^{+}J^{-}-2J^{-}|\Psi\rangle\langle\Psi|J^{+}\right)\Big{]}$
See that we obtain the corresponding second-order term of the state evolved under the Hamiltonian in Eq. (S41), $\rho_{\alpha}^{(2)}$, but we also find some new terms. In the semiclassical approximation, these are the first terms that we are neglecting. By comparing them to the largest coefficient appearing in the second order terms, which scale as $|\alpha_{P}|^{2}|\alpha_{S}|^{2}$, we can see that
[…]
$+$ $\frac{1}{|\alpha_{P}|^{2}}|\Psi\rangle\langle\Psi|J^{+}J^{-}+\frac{1}{|\alpha_{S}|^{2}}J^{-}J^{+}|\Psi\rangle\langle\Psi|+\frac{1}{|\alpha_{P}|^{2}}J^{+}J^{-}|\Psi\rangle\langle\Psi|\bigg{)}.$
Since the order of magnitude of $\alpha_{P},\ \alpha_{S}$ is $\sim 10^{7}$, these terms introduce a factor of $\sim 10^{-14}$ with respect to the leading terms, $\rho_{\alpha}^{(2)}$, and are therefore negligible. Then, the evolution of the molecular state is well-described by the propagator associated to the Hamiltonian $H_{\alpha}^{I}$ in Eq. (S41).
@@ IV Spin Coherent State
Notice that, in resonance conditions, $\Omega-\omega_{P}+\omega_{S}=0$, and there is no explicit time dependence in the Hamiltonian. Finally, we perform the semiclassical approximation described on Sec.III on the pump and Stokes modes $a_{l}\rightarrow\alpha_{l}$, and obtain
@@ V Concurrence
After vibrational molecular coherence has been generated in the molecule ensemble via stimulated Raman scattering, we will consider that one frequency of an entangled state has been sent through the fiber to change its frequency. Unlike usual stimulated Raman scattering, in which the pump beam that is scattered to the different Raman frequency sidebands needs to have a power higher than a minimum power threshold, this is a thresholdless process which can be applied to single photons. We consider $(|0,0,0\rangle+|1,1,0\rangle)/\sqrt{2}$ as the initial state, where the states in tensor product, from left to right, indicate the photon number population of the idler frequency, kept outside of the fiber, the mixing frequency, which goes initially through the fiber, and the converted frequency, initially in a vacuum state. While the phase-difference between the mixing and the converted signals has to be equal to the phase of the molecular coherence wave, the jump in frequency is given by the Raman shift. Therefore, even though this is not, properly speaking, a Raman process, we can use the same interaction Hamiltonian to describe it. Furthermore, we will focus on the interaction between the pump and the anti-Stokes, since we are looking at frequency up-conversion. Then, we start from the Hamiltonian
@@
Under phase matching and resonance conditions, we have $\beta_{U}-\beta_{M}=\Delta\beta\equiv\beta_{P}-\beta_{S}$ and $\omega_{U}-\omega_{M}=\Omega\equiv\omega_{P}-\omega_{S}$, respectively. Note that we do not require $\omega_{l}=\omega_{m}$ or $\beta_{l}=\beta_{m}$ with $l\in\{M,U\}$ and $m\in\{P,S\}$. Assuming these phase matching and resonance conditions are satisfied, this Hamiltonian simplifies to
@@
To test the validity of the semiclassical approximation on the global spin operators, we first compare the molecular states $|s\rangle$ and $|s^{\prime}\rangle$, which describe spin coherent states that differ in a single molecular excitation, such that $|s^{\prime}\rangle=J^{+}|s\rangle/\sqrt{\langle s|J^{-}J^{+}|s\rangle}$. We use the trace distance to compare these states, finding
$T\left(|s\rangle,|s^{\prime}\rangle\right)=\frac{1}{\sqrt{1+N|s|^{2}}}.$ (S84)
One can see that, as the number of molecules increases, this distance goes to zero. This suggests that, in the single-excitation subspace, the molecular state will not change significantly, provided that the number of molecules is large ($\sim 10^{18}$ in Ref. SM_Tyumenev2022, ). To go a bit further, we would like to compare the Hamiltonian in Eq. (S83) to $H_{\xi^{\prime}}^{I}$, with $\xi^{\prime}=\langle s^{\prime}|J^{-}|s^{\prime}\rangle$, in order to comprehend the magnitude of the difference between the two in the single-excitation subspace. The norm of this difference is
$\left|\left|H_{\xi}^{I}-H_{\xi^{\prime}}^{I}\right|\right|=\sqrt{2}\hbar G_{U}\left|\xi-\xi^{\prime}\right|=\sqrt{2}\hbar G_{U}|s|\left|\frac{2}{1+|s|^{2}}-\frac{N}{1+N|s|^{2}}\right|.$ (S85)
In the limit $N\rightarrow\infty$, we find
$\lim_{N\rightarrow\infty}\left|\left|H_{\xi}^{I}-H_{\xi^{\prime}}^{I}\right|\right|=\frac{\sqrt{2}\hbar G_{U}}{|s|}\left|\frac{1-|s|^{2}}{1+|s|^{2}}\right|,$ (S86)
which, as we can see at Fig. S2, is very close to zero with the parametric values used in this work. In view of these results, we conclude that the Hamiltonian in Eq. (S83) is adequate for working in the single excitation subspace.
Figure S2: Plot of the norm of the Hamiltonian difference $||H_{\xi}^{I}-H_{\xi^{\prime}}^{I}||$ in logarithmic scale in the limit $N\rightarrow\infty$ as a function of the propagation distance $z$. The value of $|s|$ is computed with the parameters used for Fig. 3 of the main text.
Now, we need to obtain the action of this Hamiltonian onto the initial photonic state of the system, which will be $(|0,0,0\rangle+|1,1,0\rangle)/\sqrt{2}$. But first, we need to transform this state into the interaction picture,
@@
We obtain the idler-mixing density matrix by doing the partial trace of $\rho$ with respect to the up-converted subsystem,
@@
and the idler-up-converted density matrix is obtained similarly from $\rho$ by tracing out the mixing frequency subsystem,
@@
$C=\max\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\},$ (S99)
@@
and hence, the concurrence is given by $C_{I-M}=\max\{0,2|w|\}$. In the case of $\rho_{I,U}$, we have
@@
finding the concurrence $C_{I-U}=\max\{0,2|y|\}$. Therefore, the concurrences can be computed as
$C_{I,M}$ $=$ $\Big{|}\cos G_{U}|\xi|t\Big{|},$ (S108)
$C_{I,U}$ $=$ $\Big{|}\sin G_{U}|\xi|t\Big{|}.$ (S109)
@@
$\mathcal{E}_{\rho}=1-\frac{1}{2}\left[\left(1+\sqrt{1-C^{2}}\right)\log_{2}\left(1+\sqrt{1-C^{2}}\right)+\left(1-\sqrt{1-C^{2}}\right)\log_{2}\left(1-\sqrt{1-C^{2}}\right)\right].$ (S110)
@@ VI Concurrence without semiclassical approximation for spins
In this section, we will look into the semiclassical approximation for spin operators that we performed before computing the concurrence between the idler-mixing and the idler-up-converted frequency modes. This approximation is based on the same one for bosonic states, in which the system is assumed to be in a coherent state $|\alpha\rangle$, and thus we replace $a\rightarrow\alpha$. This assumes that the system remains in a coherent state, which is what we are considering here with the molecules; after these are in a coherent state, we are looking at the dynamics after introducing a single-excitation into the system. Let us derive here the idler-mixing and idler-up-converted states without the semiclassical approximation of the spin operators. We start from the state
@@
in the Schrödinger picture, which describes the idler, mixing, up-converted, and molecules, in that order. We want to obtain the evolution of this state under our Hamiltonian
@@
where we have already implemented the phase-matching condition $\beta_{U}-\beta_{M}=\Delta\beta$. Again, we now move to an interaction picture with respect to $\hbar\Omega J_{z}+\hbar\omega_{M}a_{M}^{\dagger}a_{M}+\hbar\omega_{U}a_{U}^{\dagger}a_{U}$, where we have
@@
The density matrix that we obtain after performing a partial trace operation with respect to the molecular subsystem is
@@
By tracing the up-converted subsystem, we obtain the idler-mixing density matrix,
@@
while the density matrix for the idler and the up-converted frequency modes is obtained by tracing the mixing subsystem,
@@
$x$ $=$ $\sum_{n=0}^{N}|c_{1}(n)|^{2}=\frac{1}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N}\begin{pmatrix}N\\ n\end{pmatrix}|s|^{2n}\cos^{2}(G_{U}t\sqrt{n(N-n+1)}),$ (S128)
$w$ $=$ $\sum_{n=0}^{N}c_{1}(n)c_{0}^{*}(n)=\frac{1}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N}\begin{pmatrix}N\\ n\end{pmatrix}|s|^{2n}\cos(G_{U}t\sqrt{n(N-n+1)}),$ (S129)
$y$ $=$ $\sum_{n=0}^{N-1}c_{2}(n+1)c_{0}^{*}(n)=-ie^{-i(\Omega t+\Delta\beta z)}\frac{s}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N-1}\begin{pmatrix}N\\ n\end{pmatrix}\sqrt{\frac{N-n}{n+1}}|s|^{2n}\sin(G_{U}t\sqrt{(n+1)(N-n)}).$ (S130)
@@
$w=\frac{1}{2}(1+|s|^{2})^{-N}\sum_{n=0}^{N}\begin{pmatrix}N\\ n\end{pmatrix}|s|^{2n}\cos(G_{U}t\sqrt{n(N-n+1)}),$ (S131)
$w_{\text{sc}}=\frac{1}{2}\cos G_{U}|\xi|t.$ (S132)
@@
$\frac{G_{U}t}{8}\Big{[}-NG_{U}t\cos^{2}(2G_{S}\alpha_{P}\alpha_{S}t)\cos G_{U}|\xi|t+\left[\cot(2G_{S}\alpha_{P}\alpha_{S}t)-\tan(2G_{S}\alpha_{P}\alpha_{S}t)\right]\sin G_{U}|\xi|t\Big{]}.$ (S137)
@@
$\frac{1}{8N}\left[-G_{U}^{2}t^{2}|\xi|^{2}\left(1-\cot^{2}(2G_{S}\alpha_{P}\alpha_{S}t)\right)^{2}\cos G_{U}|\xi|t+G_{U}t|\xi|\left(\frac{1}{\sin^{2}(2G_{S}\alpha_{P}\alpha_{S}t)}-\frac{3}{\cos^{2}(2G_{S}\alpha_{P}\alpha_{S}t)}\right)\sin G_{U}|\xi|t\right].$ (S138)
@@
Figure S3: The next order in the idler-mixing concurrence summation plotted as a function of the propagation distance $z$ (see Eq. (S139)). The parameters for this simulation are the same ones used for the figures in the main text.
The next-order correction to the idler-mixing concurrence is
$C_{I,M}^{\text{next order}}=\left|\frac{G_{U}t}{4}\Big{[}-NG_{U}t\cos^{2}(2G_{S}\alpha_{P}\alpha_{S}t)\cos G_{U}|\xi|t+\left[\cot(2G_{S}\alpha_{P}\alpha_{S}t)-\tan(2G_{S}\alpha_{P}\alpha_{S}t)\right]\sin G_{U}|\xi|t\Big{]}\right|.$ (S139)
This is represented in Fig. S3 against the length of the fiber, where we can observe that these corrections are very small, reaching maximum values of $\sim 2\times 10^{-16}$ during the length of the fiber.
VII Equations of motion
Going back to the single-molecule case, here is shown that the Maxwell-Bloch equations describing stimulated Raman scattering SM_Raymer1990 can be recovered through the equations of motion associated to the effective Hamiltonian derived in the previous section. In the following, we only describe the pump-Stokes interaction, considering that the anti-Stokes population is negligible throughout the process. We will work with the Hamiltonian
We first start in the Schrödinger picture, and go to an interaction picture with respect to the pump and Stokes fields. We transform $H_{\text{eff}}$ by $\omega_{P}a^{\dagger}_{P}a_{P}+\omega_{S}a^{\dagger}_{S}a_{S}$, obtaining the interaction picture Hamiltonian
[…]
using the formulas
$[\sigma_{z},\sigma^{+}]$ $=$ $2\sigma^{+},$
$[\sigma_{z},\sigma^{-}]$ $=$ $-2\sigma^{-},$
$[\sigma^{+},\sigma^{-}]$ $=$ $\sigma_{z}.$
By defining $w=\rho_{1,1}-\rho_{0,0}$, we write the following EOM,
where we have not written the equation for $\dot{\rho}_{10}$ because it is just the complex conjugate of that for $\dot{\rho}_{0,1}$. The differences we find with the previous results lie in the definition of the electric field. While we have defined $\mathcal{E}(z,t)=\mathcal{E}_{P}(z,t)+\mathcal{E}_{S}(z,t)$, previous works defined $\mathcal{E}(z,t)=(\mathcal{E}_{P}(z,t)+\mathcal{E}_{S}(z,t))/2$, and thus the extra factor of $1/4$. Furthermore, while we have chosen the plane-wave solution of the electric fields as
previous works have chosen
and this is why we find $E_{P}^{*}E_{S}$ instead of $E_{P}E_{S}^{*}$.